# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ Abstract
Mediation analysis is a crucial tool for uncovering the mechanisms through which a treatment affects the outcome, providing deeper causal insights and guiding effective interventions. Despite advances in analyzing the mediation effect with fixed/low-dimensional mediators and covariates, our understanding of estimation and inference of mediation functional in the presence of (ultra)-high-dimensional mediators and covariates is still limited. In this paper, we present an estimator for mediation functional in a high-dimensional setting that accommodates the interaction between covariates and treatment in generating mediators, as well as interactions between both covariates and treatment and mediators and treatment in generating the response. We demonstrate that our estimator is $\sqrt{n}$-consistent and asymptotically normal, thus enabling reliable inference on direct and indirect treatment effects with asymptotically valid confidence intervals. A key technical contribution of our work is to develop a multi-step debiasing technique, which may also be valuable in other statistical settings with similar structural complexities where accurate estimation depends on debiasing. We evaluate our proposed methodology through extensive simulation studies and apply it to the TCGA lung cancer dataset to estimate the effect of smoking, mediated by DNA methylation, on the survival time of lung cancer patients.
Corresponding authors.
@@ 1 Introduction
The advancement of technology aids in the collection of an extensive array of variables in various disciplines such as brain imaging, genetics, epidemiology, and public health studies. Consequently, many scientific domains are faced with scenarios in which mediation analysis must be carried out in a high-dimensional setting; that is, the number of potential mediators is quite large and can even be greater than the number of observations. This challenge is particularly pronounced in environmental epigenetics when using omics data. However, most existing methodologies rely on fixed-dimensional mediators, which cannot be applied directly to high-dimensional data. This has motivated recent work on developing methodologies for high-dimensional mediation analysis. Several methodologies have been developed for estimating direct and indirect treatment effects in the presence of high-dimensional mediators. Zhang et al., (2016) and Perera et al., (2022) proposed to first screen mediators that are strongly associated with outcome and run standard analysis on these chosen low-dimensional mediators. Huang and Pan, (2016) and Chén et al., (2018) suggested reducing dimension by a general mapping to low-dimensional spaces instead of variable selection. Zhang et al., (2018) also performed dimension reduction but based on using distance metrics and grouping individually weak signals. Zhao and Luo, (2016) proposed a pathwise Lasso method for selecting sparse mediation pathways in the presence of high-dimensional mediators. None of the aforementioned works presented any theoretical study for their proposed methodology. However, limited theoretical progress has also been made in the literature. For example, Guo et al., (2022) provides an estimator (along with theoretical guarantees) of direct and indirect treatment effects in the presence of high-dimensional mediators and fixed-dimensional observed confounders and exposures. Guo et al., (2023) considered testing the effect the high-dimensional mediators on the response variable in the presence of fixed dimensional exposures. Guo et al., (2024) expanded the estimation method of Guo et al., (2023) by incorporating a general link function to characterize the relationship between the response variable and the mediators and exposures. Lin et al., (2023) proposed a methodology to estimate the indirect effect in the presence of multivariate mediators and exposures. Their method allows both the dimension of the mediators and the exposures to grow with the sample size. A recent study by Rakshit and Guo, (2024) introduced an estimator for the direct (and indirect) treatment effect, accommodating high-dimensional mediators and their interaction with treatment. Our approach generalizes this work within a linear model framework, extending it to include both high-dimensional covariates, mediators, and their interaction with the treatment indicator. We note that all these papers either consider no observed covariates or fixed-dimensional observed covariates in the setting, i.e., they do not allow both the mediators and the observed covariates to be (ultra) high-dimensional. Moreover, they do not allow for interaction between exposures and mediators or between exposures and the observed covariates.
@@
Our contribution: In this paper, we propose a novel methodology for estimating the mediation functional in the presence of ultra-high-dimensional (UHD) covariates and mediators (where the dimension can be exponentially large compared to the sample size). We consider the case where the response variable is generated from a UHD linear model (with respect to both the covariates and the mediators), and the UHD mediators themselves are generated from a linear model with respect to the UHD covariates. Unlike the majority of the existing literature on high-dimensional mediation analysis, our setup allows for interaction with binary exposure for both data-generating processes. To the best of our knowledge, this is the first work to provide a methodology for estimating direct and indirect treatment effects in the presence of both UHD covariates and mediators, along with their interaction with a binary exposure variable. Our method of estimating mediation functional essentially consists of three high-dimensional regressions. We first regress response on both covariates and mediators using treatment observations and then regress mediators on covariates using control observations. Since both regressions are high-dimensional, we carefully use a debiasing technique to remove the bias, which necessitates the third high-dimensional regression. We demonstrate that, under fairly general assumptions, our proposed estimator of mediation functional is $\sqrt{n}$-consistent and asymptotically normal (where $n$ is the sample size). We also provide a consistent estimator for the asymptotic variance, which aids in constructing an asymptotically valid confidence interval.
@@
Our debiasing approach is primarily motivated by the large body of literature on the debiasing techniques for estimating a contrast of high-dimensional parameters in a linear model (see, e.g., (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)). The use of the debiasing technique in causal inference was also recently employed by Athey et al., (2018), where the authors developed a methodology to estimate the total effect of a binary treatment on an outcome in the presence of high-dimensional covariates. We generalize that work by estimating the direct and indirect parts of the causal effect separately while allowing the mediator to be UHD and have interactions with the treatment variable. Our proposed method can be considered as a high-dimensional counterpart of the influence function-based approach of Tchetgen Tchetgen and Shpitser, (2012). In particular, we demonstrate that our estimator possesses a second-order bias with product form, similar to the influence function-based approach for the fixed-dimensional setting, as established by Ghassami, (2024) based on the influence function of Tchetgen Tchetgen and Shpitser, (2012).
@@
The rest of the paper is organized as follows. After introducing the problem setting in Section 2, we describe our debiased estimation procedure in Section 3. In Section 4, we provide the bias analysis of our estimator and its connection to the fixed-dimensional setting. In Section 5, we provide a theoretical justification for our approach under high-dimensional asymptotic. The proof of the main theorem, along with technical lemmas, can be found in the Appendix. In Section 6, we conduct a simulation experiment and a real data experiments on TCGA Lung cancer dataset. A software implementation for R is available at UHDmedi. We conclude in Section 7 and discuss future work.
@@ 2 Notations and Problem Setting
For a random variable/vector $X$, we denote $\mathbb{E}[X]$ and $\Sigma_{X}$ as the expectation and the variance/covariance matrix of $X$, respectively. For a matrix $\mathbf{X}$, we use $\mathbf{X}_{i,.}$ (respectively $\mathbf{X}_{.,i}$) to denote its $i^{th}$ row (respectively $i^{th}$ column). For a matrix $\mathbf{X}$, $\|\mathbf{X}\|_{1}$, $\|\mathbf{X}\|_{2}$, $\|\mathbf{X}\|_{\infty}=\max_{i,j}|\mathbf{X}_{ij}|$, $\|\mathbf{X}\|_{1,\infty}=\underset{1\leq i\leq m}{\max}\|\mathbf{X}_{i,\cdot}\|_{1}$, $\|\mathbf{X}\|_{\text{op}}=\underset{\|v\|=1}{\sup}\|\mathbf{X}v\|$ and $\|\mathbf{X}\|_{F}=\sqrt{\sum_{i,j}X^{2}_{i,j}}$ represent its $L_{1}$-norm, $L_{2}$-norm, $L_{\infty}$-norm, $L_{1,\infty}$-norm, operator norm and Frobenius norm, respectively. $\mathbf{1}_{q}$ and $\mathbf{0}_{q}$ are the vectors consisting of all ones and all zeros, respectively in $\mathbb{R}^{q}$. $I_{n}$ represents the identity matrix of size $n$. $\mathcal{N}(\mu,\sigma^{2})$ represents the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. We denote maximum and minimum by $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
@@
$\mathbb{E}\left[Y^{(1)}-Y^{(0)}\right]=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(0,M^{(0)}\right)}\right]=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(1,M^{(0)}\right)}\right]+\mathbb{E}\left[Y^{\left(1,M^{(0)}\right)}-Y^{\left(0,M^{(0)}\right)}\right].$ (2.1)
The first and the second terms in the last expression are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992), and are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively by Pearl, (2001). We will use the latter terminology in this work. NIE captures the change in the expectation of the response variable in a hypothetical scenario where the value of the treatment variable is fixed at $A=1$, while the mediator behaves as if the treatment had been changed from value 0 to 1. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value 0 to 1, while the mediator behaves as if the treatment is fixed at 0. Note that if one can identify and estimate the parameter $\mathbb{E}[Y^{(a^{\prime},M^{(a)})}]$ for $a,a^{\prime}\in\{0,1\}$, then based on Equation (2.1), NIE, NDE and ATE can all be estimated. Henceforth, we focus on the parameter $\mathbb{E}[Y^{(1,M^{(0)})}]$ in this paper.
@@
    it allows for $A-X$ and $A-M$ interactions (although it requires no $X-M$ interaction).
@@
for all $X\in\mathbb{R}^{p},M\in\mathbb{R}^{q}$, where we absorbed the intercepts in the means. Therefore, the parameter of interest in Equation (2.2) reduced to
@@ 3 Debiased Estimator for the Mediation Functional
[…]
@@
We now present the details of our methodology, which is summarized in Algorithm 1 and in Figure 1. Our method essentially comprises five key steps. In the first step, we obtain $\hat{\beta}_{1},\hat{\gamma}_{1}$ by regressing $Y$ on $(X,M)$ using the treatment observations. As we are in a UHD regime, we use the $\ell_{1}$ penalty for consistent estimation. We also regress $M$ on $X$ using control observations, along with $\ell_{1}$ penalty, to obtain $\hat{\mathbf{B}}_{0}$. As $M$ is multivariate, we regress each coordinate of $M$ on $X$ in parallel to estimate the rows of $\mathbf{B}_{0}$.
@@
Step 1: Split the data into two parts, say $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Use $\mathcal{D}_{1}$ to run $q$ parallel regressions of $(\mathbf{M}_{c})_{.,j},j=1,\ldots,q$ on $\mathbf{X}_{c}$, to obtain $\hat{\mathbf{B}}_{0_{1,.}},\ldots,\hat{\mathbf{B}}_{0_{q,.}}$ where each $\hat{\mathbf{B}}_{0_{i,.}}\in{\mathbb{R}}^{p}$. Concatenate them to obtain the following estimator for $\mathbf{B}_{0}$
$\hat{\mathbf{B}}_{0}=[\hat{\mathbf{B}}_{0_{1,.}}\quad\hat{\mathbf{B}}_{0_{2,.}}\quad\cdots\quad\hat{\mathbf{B}}_{0_{q,.}}]^{\top}.$
Use $\mathcal{D}_{2}$ to Estimate ${\phi}_{1}=[{\beta}_{1}^{\top}\quad{\gamma}_{1}^{\top}]^{\top}$ by regressing $Y_{t}$ on $\mathbf{W}_{t}=[\mathbf{X}_{t}\quad\mathbf{M}_{t}]$ along with a $\ell_{1}$ penalty
$\hat{\phi}_{1}=[{\hat{\beta}}_{1}^{\top}\quad{\hat{\gamma}}_{1}^{\top}]^{\top}=\arg\min_{\tilde{\phi}_{1}}\left\{\sum_{\{i:A_{i}=1\}}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\tilde{\phi}_{1}\right)^{2}+\lambda_{1}\|\tilde{\phi}_{1}\|_{1}\right\}.$
Step 2: Compute bias correction weights $\tau_{1}$ with contrast $a=\left[\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\right]^{\top}$ and tuning parameter $K_{1}$ using $\mathcal{D}_{2}$,
@@
$\hat{\theta}_{0,1}=a^{\top}\hat{\phi}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right).$
@@
$\hat{b}=\arg\min_{\tilde{b}}\left\{\sum_{\{i:A_{i}=0\}}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\tilde{b}\right)^{2}+\lambda_{2}\|\tilde{b}\|_{1}\right\}.$
@@
In the second step, we debias $\hat{\phi}_{1}:=(\hat{\beta}_{1},\hat{\gamma}_{1})$ in the direction of $a:=(\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X})$ to obtain an intermediate estimator $\hat{\theta}_{0,1}$. For notational simplicity, we define $\mathbf{W}=(\mathbf{X}\quad\mathbf{M})$ and $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$. More specifically, we have:
$\hat{\theta}_{0,1}=\bar{X}^{\top}\hat{\beta}_{1}+(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right)\ \equiv\ \hat{\theta}^{\rm plug-in}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right)\,,$
@@
To remove this bias, in the third step, we first regress $M^{\top}\hat{\gamma}_{1}$ on $X$ along with an $\ell_{1}$ penalty (as $X$ is a UHD vector). As $\mathbb{E}[M\mid X,A=0]=\mathbf{B}_{0}X$, this penalized regression effectively estimates $\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$. In the fourth step, we debias this estimator along the direction of $\bar{X}$, i.e., we obtain
$\hat{\theta}_{0,2}=\bar{X}\cdot\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\hat{b}\right),$
where $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, and $\hat{b}=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, and the additional summand is the correction term obtained by solving a suitable optimization problem (see Equation (3.2) for details) that aims to reduce the bias of the penalized estimator. In the final step, we obtain our final debiased estimator $\hat{\theta}_{0}$ as

### v3

@@ Abstract
Mediation analysis is a crucial tool for uncovering the mechanisms through which a treatment affects the outcome, providing deeper causal insights and guiding effective interventions. Despite advances in analyzing the mediation effect with fixed/low-dimensional mediators and covariates, our understanding of estimation and inference of mediation functional in the presence of (ultra)-high-dimensional mediators and covariates is still limited. In this paper, we present an estimator for mediation functional in a high-dimensional setting that accommodates the interaction between covariates and treatment in generating mediators, as well as interactions between both covariates and treatment and mediators and treatment in generating the response. We demonstrate that our estimator is $\sqrt{n}$-consistent and asymptotically normal, thus enabling reliable inference on direct and indirect treatment effects with asymptotically valid confidence intervals. A key technical contribution of our work is to develop a multi-step debiasing technique, which may also be valuable in other statistical settings with similar structural complexities where accurate estimation depends on debiasing. We evaluate our proposed methodology through extensive simulation studies and apply it to the TCGA lung cancer dataset to estimate the effect of smoking, mediated by DNA methylation, on the survival time of lung cancer patients. Corresponding authors.
@@ 1 Introduction
The advancement of technology aids in the collection of an extensive array of variables in various disciplines such as brain imaging, genetics, epidemiology, and public health studies. Consequently, many scientific domains are faced with scenarios in which mediation analysis must be carried out in a high-dimensional setting; that is, the number of potential mediators is quite large and can even be greater than the number of observations. This challenge is particularly pronounced in environmental epigenetics when using omics data. However, most existing methodologies rely on fixed-dimensional mediators, which cannot be applied directly to high-dimensional data. This has motivated recent work on developing methodologies for high-dimensional mediation analysis. Several methodologies have been developed for estimating direct and indirect treatment effects in the presence of high-dimensional mediators. Zhang et al., (2016) and Perera et al., (2022) proposed to first screen mediators that are strongly associated with the outcome and run standard analysis on these chosen low-dimensional mediators. Huang and Pan, (2016) and Chén et al., (2018) suggested reducing dimension by a general mapping to low-dimensional spaces instead of variable selection. Zhang et al., (2018) also performed dimension reduction, but based on using distance metrics and grouping individually weak signals. Zhao and Luo, (2016) proposed a pathwise Lasso method for selecting sparse mediation pathways in the presence of high-dimensional mediators. None of the aforementioned works presented any theoretical study for their proposed methodology. However, limited theoretical progress has also been made in the literature. For example, Guo et al., (2022) provides an estimator (along with theoretical guarantees) of direct and indirect treatment effects in the presence of high-dimensional mediators and fixed-dimensional observed confounders and exposures. Guo et al., (2023) considered testing the effect of the high-dimensional mediators on the response variable in the presence of fixed dimensional exposures. Guo et al., (2024) expanded the estimation method of Guo et al., (2023) by incorporating a general link function to characterize the relationship between the response variable and the mediators and exposures. Lin et al., (2023) proposed a methodology to estimate the indirect effect in the presence of multivariate mediators and exposures. Their method allows both the dimension of the mediators and the exposures to grow with the sample size. A recent study by Rakshit and Guo, (2024) introduced an estimator for the direct (and indirect) treatment effect, accommodating high-dimensional mediators and their interaction with treatment. Our approach generalizes this work within a linear model framework, extending it to include both high-dimensional covariates, mediators, and their interaction with the treatment indicator. We note that all these papers either consider no observed covariates or fixed-dimensional observed covariates in the setting, i.e., they do not allow both the mediators and the observed covariates to be (ultra) high-dimensional. Moreover, they do not allow for interaction between exposures and mediators or between exposures and the observed covariates.
@@
Our contribution: In this paper, we propose a novel methodology for estimating the mediation functional in the presence of ultra-high-dimensional (UHD) covariates and mediators (where the dimension can be exponentially large compared to the sample size). We consider the case where the response variable is generated from a UHD linear model (with respect to both the covariates and the mediators), and the UHD mediators themselves are generated from a linear model with respect to the UHD covariates. Unlike the majority of the existing literature on high-dimensional mediation analysis, our setup allows for interaction with binary exposure for both data-generating processes. To the best of our knowledge, this is the first work to provide a methodology for estimating direct and indirect treatment effects in the presence of both UHD covariates and mediators, along with their interaction with a binary exposure variable. Our method of estimating mediation functional essentially consists of three high-dimensional regressions. We first regress response on both covariates and mediators using treatment observations and then regress mediators on covariates using control observations. Since both regressions are high-dimensional, we carefully use a debiasing technique to remove the bias, which necessitates the third high-dimensional regression. We demonstrate that, under fairly general assumptions, our proposed estimator of the mediation functional is $\sqrt{n}$-consistent and asymptotically normal (where $n$ is the sample size). We also provide a consistent estimator for the asymptotic variance, which aids in constructing an asymptotically valid confidence interval.
@@
Our debiasing approach is primarily motivated by the large body of literature on the debiasing techniques for estimating a contrast of high-dimensional parameters in a linear model (see, e.g., (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)). The use of the debiasing technique in causal inference was also recently employed by Athey et al., (2018), where the authors developed a methodology to estimate the total effect of a binary treatment on an outcome in the presence of high-dimensional covariates. We generalize that work by estimating the direct and indirect parts of the causal effect separately while allowing the mediator to be UHD and have interactions with the treatment variable. Our proposed method can be considered as a high-dimensional counterpart of the influence function-based approach of Tchetgen Tchetgen and Shpitser, (2012). In particular, we demonstrate that our estimator possesses a second-order bias with product form, similar to the influence function-based approach for the fixed-dimensional setting, as established by Liu and Ghassami, (2024) based on the influence function of Tchetgen Tchetgen and Shpitser, (2012).
@@
The rest of the paper is organized as follows. After introducing the problem setting in Section 2, we describe our debiased estimation procedure in Section 3. In Section 4, we provide the bias analysis of our estimator and its connection to the fixed-dimensional setting. In Section 5, we provide a theoretical justification for our approach under high-dimensional asymptotics. The proof of the main theorem, along with technical lemmas, can be found in the supplementary material. In Section 6, we conduct some simulation experiments and a real data experiment on the TCGA Lung cancer dataset. A software implementation for R is available at UHDmedi. We conclude in Section 7 and discuss future work.
@@ 2 Notations and Problem Setting
For a random variable/vector $X$, we denote $\mathbb{E}[X]$ and $\Sigma_{X}$ as the expectation and the variance/covariance matrix of $X$, respectively. For a matrix $\mathbf{X}$, we use $X_{i}$ (respectively $X_{.,i}$) to denote its $i^{th}$ row (respectively $i^{th}$ column). For a matrix $\mathbf{X}$, $\|\mathbf{X}\|_{1}$, $\|\mathbf{X}\|_{2}$, $\|\mathbf{X}\|_{\infty}=\max_{i,j}|X_{ij}|$, $\|\mathbf{X}\|_{1,\infty}=\underset{1\leq i\leq m}{\max}\|X_{i,\cdot}\|_{1}$, $\|\mathbf{X}\|_{\text{op}}=\underset{\|v\|=1}{\sup}\|\mathbf{X}v\|$ and $\|\mathbf{X}\|_{\text{F}}=\sqrt{\sum_{i,j}X^{2}_{i,j}}$ represent its $L_{1}$-norm, $L_{2}$-norm, $L_{\infty}$-norm, $L_{1,\infty}$-norm, operator norm and Frobenius norm, respectively. $\mathbf{1}_{q}$ and $\mathbf{0}_{q}$ are the vectors consisting of all ones and all zeros, respectively, in $\mathbb{R}^{q}$. $I_{n}$ represents the identity matrix of size $n$. $\mathcal{N}(\mu,\sigma^{2})$ represents the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. We denote maximum and minimum by $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
@@
$\mathbb{E}\left[Y^{(1)}-Y^{(0)}\right]$ $=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(0,M^{(0)}\right)}\right]$ (2.1)
$=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(1,M^{(0)}\right)}\right]+\mathbb{E}\left[Y^{\left(1,M^{(0)}\right)}-Y^{\left(0,M^{(0)}\right)}\right].$
The first and the second terms in the last expression are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992), and are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively, by Pearl, (2001). We will use the latter terminology in this work. NIE captures the change in the expectation of the response variable in a hypothetical scenario where the value of the treatment variable is fixed at $A=1$, while the mediator behaves as if the treatment had been changed from value 0 to 1. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value 0 to 1, while the mediator behaves as if the treatment is fixed at 0. Note that if one can identify and estimate the parameter $\mathbb{E}[Y^{(a^{\prime},M^{(a)})}]$ for $a,a^{\prime}\in\{0,1\}$, then based on Equation (2.1), NIE, NDE and ATE can all be estimated. Henceforth, we focus on the parameter $\mathbb{E}[Y^{(1,M^{(0)})}]$ in this paper.
@@
    it allows for AXA-X and AMA-M interactions (although it requires no XMX-M interaction).
@@
for all $X\in\mathbb{R}^{p},M\in\mathbb{R}^{q}$, where we absorbed the intercepts in the means. Therefore, the parameter of interest in Equation (2.2) is reduced to
@@ 3 Debiased Estimator for the Mediation Functional
$\hat{\beta}_{1}$$\hat{\gamma}_{1}$$\widehat{\mathbf{B_{0}^{\top}\widehat{\gamma}}_{1}}$$\hat{\theta}_{0,1}$$\hat{\theta}_{0,2}$$\hat{\theta}_{0}$$\hat{\mathbf{B}}_{0}$$\begin{array}[]{l}Regress\ Y\ on\ (X,\ M)\\ \ \ \ \ \ \ \ (on\ treated)\end{array}$$\begin{array}[]{l}Regress\ M_{i}^{\top}\hat{\gamma}_{1}\ on\ X_{i}\\ \ \ \ \ \ \ \ (on\ control)\end{array}$$\begin{array}[]{l}Regress\ parallely\ M\ on\ X\\ \ \ \ \ \ \ \ \ \ \ \ (on\ control)\end{array}$$\begin{array}[]{l}Debias\ (\hat{\beta}_{1},\ \hat{\gamma}_{1})\ in\ the\\ direction\ (\overline{X},\ \hat{\mathbf{B}}_{0}\overline{X})\end{array}$$\begin{array}[]{l}Debias\ in\ the\\ direction\ \overline{X}\end{array}$$\hat{\theta}_{0,1}\ +\ \hat{\theta}_{0,2}\ -\ (\hat{\mathbf{B}}_{0}\overline{X})^{\top}\hat{\gamma}_{1}$
@@
We now present the details of our methodology, which is summarized in Algorithm 1 and in Figure 1. Our method essentially comprises five key steps. In Step 1, we divide the dataset into two subsets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ for estimation purposes. On the first subset, we regress $M$ on $X$ using control observations, along with $\ell_{1}$ penalty, to obtain $\hat{\mathbf{B}}_{0}$. As $M$ is multivariate, we regress each coordinate of $M$ on $X$ in parallel to estimate the rows of $\mathbf{B}_{0}$. The rest of the operations in the algorithm are performed on $\mathcal{D}_{2}$. We first use this subset of the data to obtain the estimates $\hat{\beta}_{1},\hat{\gamma}_{1}$ by regressing $Y$ on $(X,M)$ using the treatment observations. As we are in a UHD regime, we use the $\ell_{1}$ penalty for consistent estimation in both the regressions.
@@
Step 1: Split the data into two parts, say $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Use $\mathcal{D}_{1}$ to run $q$ parallel regressions of $(M_{c})_{.,j},j=1,\ldots,q$ on $\mathbf{X}_{c}$ along with a $\ell_{1}$ penalty
$\hat{B}_{0_{k}}=\operatorname{argmin}_{\tilde{B}_{0_{k}}}\left\{\sum_{\{i:A_{i}=0\}}\left((M_{c})_{.,j}-X_{i}^{\top}\tilde{B}_{0_{k}}\right)^{2}+\lambda_{0}\|\tilde{B}_{0_{k}}\|_{1}\right\}.$
to obtain $\hat{B}_{0_{1}},\ldots,\hat{B}_{0_{q}}$ where each $\hat{B}_{0_{k}}\in\mathbb{R}^{p}$. Concatenate them to obtain the following estimator for $\mathbf{B}_{0}$
$\hat{\mathbf{B}}_{0}=[\hat{B}_{0_{1}}\quad\hat{B}_{0_{2}}\quad\cdots\quad\hat{B}_{0_{q}}]^{\top}.$
Use $\mathcal{D}_{2}$ to Estimate ${\phi}_{1}=({\beta}_{1}^{\top}\quad{\gamma}_{1}^{\top})^{\top}$ by regressing $Y_{t}$ on $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$ along with a $\ell_{1}$ penalty
$\hat{\phi}_{1}=({\hat{\beta}}_{1}^{\top}\quad{\hat{\gamma}}_{1}^{\top})^{\top}=\operatorname{argmin}_{\tilde{\phi}_{1}}\left\{\sum_{\{i:A_{i}=1\}}\left(Y_{i}-W_{i}^{\top}\tilde{\phi}_{1}\right)^{2}+\lambda_{1}\|\tilde{\phi}_{1}\|_{1}\right\}.$
Step 2: Compute bias correction weights $\tau_{1}$ with contrast $a=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ and tuning parameter $K_{1}$ using $\mathcal{D}_{2}$,
@@
$\hat{\theta}_{0,1}=a^{\top}\hat{\phi}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-W_{i}^{\top}\hat{\phi}_{1}\right).$
@@
$\hat{b}=\operatorname{argmin}_{\tilde{b}}\left\{\sum_{\{i:A_{i}=0\}}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\tilde{b}\right)^{2}+\lambda_{2}\|\tilde{b}\|_{1}\right\}.$
@@
In Step 2, we debias $\hat{\phi}_{1}:=(\hat{\beta}_{1}^{\top}\quad\hat{\gamma}_{1}^{\top})^{\top}$ in the direction of $a=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ to obtain an intermediate estimator $\hat{\theta}_{0,1}$. For notational simplicity, we define $\mathbf{W}=(\mathbf{X}\quad\mathbf{M})$ and $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$. More specifically, we have:
$\hat{\theta}_{0,1}$ $=\bar{X}^{\top}\hat{\beta}_{1}+(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-W_{i}^{\top}\hat{\phi}_{1}\right)$
$\equiv\ \hat{\theta}^{\rm plug-in}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-W_{i}^{\top}\hat{\phi}_{1}\right)\,$,
@@
To remove this bias, in Step 3, we first regress $M^{\top}\hat{\gamma}_{1}$ on $X$ along with an $\ell_{1}$ penalty (as $X$ is a UHD vector). As $\mathbb{E}[M\mid X,A=0]=\mathbf{B}_{0}X$, this penalized regression effectively estimates $\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$. In Step 4, we debias this estimator along the direction of $\bar{X}$, i.e., we obtain
[…]
where $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, and $\hat{b}=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, and the additional summand is the correction term obtained by solving a suitable optimization problem (see Equation (3.2) for details) that aims to reduce the bias of the penalized estimator. In Step 5, we obtain our final debiased estimator $\hat{\theta}_{0}$ as
@@
Remark 3.1 (Cross-fitting).
In our proposed methodology, we use sample splitting, which aids in the theoretical analysis of our estimator, albeit at the expense of reduced efficiency. However, to gain efficiency, we can perform cross-fitting (Chernozhukov et al., , 2018), i.e., we reverse the role of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ and repeat the steps, and finally take the average of these two estimators.